eviant:

PLS HELP!!! Find the distance between R(–1, 1) and S(–4, –5) to the nearest tenth.

eviant:

6.7 9.0 15.8 45.0

eviant:

@Vocaloid I came up w/ 1.5, but that's not an option

Mehek:

Did you use the distance formula?

eviant:

yes

Mehek:

Do you have your work? Maybe I can see where you went wrong

eviant:

so, I did: -1-4/2 which is -2.5

eviant:

then I did -1+5/2 which is 4

Mehek:

I think you did the midpoint formula not distance

eviant:

I added those both and came up with 1.5

Mehek:

Distance formula is \(\bf\sqrt{(y_2-y_1)^2+(x_2-x_1)^2}\)

eviant:

oh

Mehek:

So let's plug in the values x1 would be -1, x2 would be -4, y1 would be 1, and y2 would be -5

Mehek:

Plugging them: \(\bf\sqrt{(-5-1)^2+(-4--1)^2}\) Could you solve that

eviant:

I'm coming with 7.8 as the final result?

eviant:

@Mehek

Mehek:

I got something different

Mehek:

Ok let's do this step by step -5 - 1 = ?